(1) How much amount (in milligrams) of riboflavin (MW=376.4 g/mol) you will need to prepare a 6.4x10-4M stock solution in 10 ml volume?

(2) How much volume of the above stock solution will you need to prepare the riboflavin solutions of following concentrations:

a. 0.059 mM in 25 ml

b. 14 uM in 25 ml
Answer #1

1) Molarity = (W/MW) (1000 / volume of solution in mL)

0.00064 = (W / 376.4) (1000 / 10)

0.0000064 = W / 376.4

W = 0.002409 g = 2.409 mg

answer = 2.409 mg

2) a) according to dilution law

M1V1 = M2V2

M1 = 0.00064 , V1 = ?

M2 = 0.059 mM = 0.000059M , V2 = 25mL

V1 = (0.000059 x 25 / 0.00064)

V1 = 2.305 mL solution required

b) M1 = 0.00064 M , V1 = ?

M2 = 14 microM = 0.000014 M , V2 = 25 mL

V1 = (0.000014 x 25 / 0.00064)

V1 = 0.547 mL solution required